# BaB-ND: Long-Horizon Motion Planning with Branch-and-Bound and Neural Dynamics

### v2

@@ C.5 Performance Change with Varying Input Discontinuities
We conducted a follow-up experiment by removing the obstacles (non-feasible regions) in the problem of Pushing with Obstacles, simplifying the objective function. Below, we report the performance of different methods on the simplified objective function (w/o obstacles) and the original objective function (w/ obstacles) in Table 7.
@@ D.2 Data Collection
For training the dynamics model, we randomly collect interaction data from simulators. For Pushing with Obstacles, Object Merging, and Object Sorting tasks, we use Pymunk (Blomqvist, 2022) to collect data, and for the Rope Routing task, we use NVIDIA FleX (Macklin et al., 2014) to generate data. In the following paragraphs, we introduce the data generation process for different tasks in detail.
@@ D.3 Details of Neural Dynamics Model Learning
We learn the neural dynamics model from the state-action pairs collected from interactions with the environment. Let the state and action at time $t$ be denoted as $x_{t}$ and $u_{t}$. Our goal is to learn a predictive model $f_{dyn}$, instantiated as a neural network, that takes a short sequence of states and actions with $l$-step history and predicts the next state at time $t+1$:
$\hat{x}_{t+1}=f_{dyn}(x_{t},u_{t}).$ (27)
@@
$\mathcal{L}=\frac{1}{T_{h}}\sum_{t=l+1}^{l+T_{h}}\|x_{t+1}-f_{dyn}(\hat{x}_{t},u_{t})\|^{2}_{2}.$ (28)
For different tasks, we choose different types of model architecture and design different inputs and outputs. For Pushing with Obstacles, Object Merging, and Rope Routing tasks, we use an MLP as our dynamics model. For the Object Sorting task, we utilize a GNN as the dynamics model, since the pieces are naturally modeled as a graph. Below is the detailed information for each task.
@@ D.5 Details of Real World Deployment
For perception, we filter the point clouds based on color from four cameras and use ICP alignment with the provided object mesh to determine the object states.
@@
For perception, we begin by using GroundingDINO and SAM to generate the mask for the rope and extract its corresponding point cloud. Subsequently, we apply farthest point sampling to identify 10 key points on the rope, representing its object state.
@@
For perception, we filter the point clouds based on color from four cameras and use K-means clustering to separate different object pieces.
@@ D.7 Implementation Details of Conventional Motion Planning Approaches
7:     Sample 1000 actions within $\{\overline{u},\underline{u}\}$ as a set $\mathcal{U}$, and compute the corresponding next states $\mathcal{X}_{\text{new}}=f_{dyn}(x_{\text{near}},\mathcal{U})$

### v3

@@ C.5 Performance Change with Varying Input Discontinuities
We conducted a follow-up experiment by removing the obstacles (non-feasible regions) in the problem of Pushing w/ Obstacles, simplifying the objective function. Below, we report the performance of different methods on the simplified objective function (w/o obstacles) and the original objective function (w/ obstacles) in Table 7.
@@ D.2 Data Collection
For training the dynamics model, we randomly collect interaction data from simulators. For Pushing w/ Obstacles, Object Merging, and Object Sorting tasks, we use Pymunk (Blomqvist, 2022) to collect data, and for the Rope Routing task, we use NVIDIA FleX (Macklin et al., 2014) to generate data. In the following paragraphs, we introduce the data generation process for different tasks in detail.
@@ D.3 Details of Neural Dynamics Model Learning
We learn the neural dynamics model from the state-action pairs collected from interactions with the environment. Let the state and action at time $t$ be denoted as $x_{t}$ and $u_{t}$. Our goal is to learn a predictive model $f_{\text{dyn}}$, instantiated as a neural network, that takes a short sequence of states and actions with $l$-step history and predicts the next state at time $t+1$:
$\hat{x}_{t+1}=f_{\text{dyn}}(x_{t},u_{t}).$ (27)
@@
$\mathcal{L}=\frac{1}{T_{h}}\sum_{t=l+1}^{l+T_{h}}\|x_{t+1}-f_{\text{dyn}}(\hat{x}_{t},u_{t})\|^{2}_{2}.$ (28)
For different tasks, we choose different types of model architecture and design different inputs and outputs. For Pushing w/ Obstacles, Object Merging, and Rope Routing tasks, we use an MLP as our dynamics model. For the Object Sorting task, we utilize a GNN as the dynamics model, since the pieces are naturally modeled as a graph. Below is the detailed information for each task.
@@ D.5 Details of Real World Deployment
For perception, we filter the point cloud based on color from four cameras, and align it with iterative closest point (ICP) to determine the object states.
@@
For perception, we begin by using Grounding DINO (Liu et al., 2023a) and Segment Anything Model (SAM) (Kirillov et al., 2023) to generate the mask for the rope and extract its corresponding point cloud. Subsequently, we apply farthest point sampling to identify 10 key points on the rope, representing its object state.
@@
For perception, we filter the point cloud based on color from four cameras and use K-means clustering to separate different object pieces.
@@ D.7 Implementation Details of Conventional Motion Planning Approaches
7:     Sample 1000 actions within $\{\overline{u},\underline{u}\}$ as a set $\mathcal{U}$, and compute the corresponding next states $\mathcal{X}_{\text{new}}=f_{\text{dyn}}(x_{\text{near}},\mathcal{U})$